OpenStudy (jgresham5):

The price of a technology stock was $9.55 yesterday. Today, the price rose to $9.64. Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (scooby6363):

To find percentage increase/decrease: 1. Find the difference between the two values. 9.64 - 9.55 = .09 2. Divide the difference by the original value in the question. .09 / 9.55 = .009424 3. Multiply the result by 100 in order to find the percent increase/decrease. .009424 * 100 = .9424 This is your answer. If you get another question like this and the number is negative that means that a percentage decrease occurred.